# Robust Ladder Climbing with a Quadrupedal Robot

### v1

@@ Abstract
Quadruped robots are proliferating in industrial environments where they carry sensor suites and serve as autonomous inspection platforms. Despite the advantages of legged robots over their wheeled counterparts on rough and uneven terrain, they are still yet to be able to reliably negotiate ubiquitous features of industrial infrastructure: ladders. Inability to traverse ladders prevents quadrupeds from inspecting dangerous locations, puts humans in harm’s way, and reduces industrial site productivity. In this paper, we learn quadrupedal ladder climbing via a reinforcement learning-based control policy and a complementary hooked end-effector. We evaluate the robustness in simulation across different ladder inclinations, rung geometries, and inter-rung spacings. On hardware, we demonstrate zero-shot transfer with an overall 90% success rate at ladder angles ranging from 70 to 90, consistent climbing performance during unmodeled perturbations, and climbing speeds 232$\times$ faster than the state of the art. This work expands the scope of industrial quadruped robot applications beyond inspection on nominal terrains to challenging infrastructural features in the environment, highlighting synergies between robot morphology and control policy when performing complex skills. More information can be found at the project website: https://sites.google.com/leggedrobotics.com/climbingladders.
@@ I INTRODUCTION
Historically, most quadruped robots were limited to locomotion on highly structured terrains [1]. Advances in control architectures and hardware over the past decade have resulted in systems which can perform feats of agility, including stable locomotion over irregular natural surfaces [2, 3] and “parkour”-like leaping over obstacles [4, 5]. Nowadays, due to their ability to traverse rough terrain and remain stable on uneven surfaces, quadruped robots are becoming commonplace at industrial sites where they perform routine inspection tasks in environments that are dangerous or undesirable to humans [6]. Despite ongoing advances, quadruped robots are still unable to robustly traverse many types of infrastructure that are common in man-made environments, a key example being ladders.
@@
Figure 1: Composite image of a quadrupedal robot equipped with hooked end-effectors, ascending a ladder in 4 s with a reinforcement learning-based control policy. Ladder shown has parameters 90 incline, 1.8 m length, 0.3 m inter-rung spacing, 2.5 cm rung radius, and 1 m width.
@@
Ladder falls are a major occupational hazard; the American Academy of Orthopedic Surgeons estimates that 500,000 ladder-related injuries occur per year in the United States alone [7]. To minimize occupational risk and improve site efficiency, next-generation industrial inspection robots must be able to robustly negotiate all types of industrial terrain, including ladders. However, quadrupeds are not normally equipped with the appropriate morphology or control policy for such a task. Consider the following:
  • Quadruped legs are usually designed with ball-shaped or flat feet which preclude the generation of anchoring forces required for reliable upward movement [8].
  • The mechanics of ladder climbing involve full-body coordination to stabilize the center of mass and move upward at a high degree of inclination [9]—something that current locomotion controllers cannot accomplish reliably in unmodeled or “noisy” environments [10].
  • Ladders may vary widely in total length ($L_{len}$), width ($L_{width}$), inter-rung spacing ($L_{space}$), rung type (flat or round), rung radius ($L_{radius}$), and degree of inclination ($L_{\theta}$), requiring robust generalization of a controller.
Indeed, robotic ladder climbing has proven an elusive task, posing many interconnected challenges that have been the subjects of study for decades.
@@
Previous research on robotic ladder climbing has been conducted mainly with humanoid robots, been characterized by very slow climbing speeds, and has not generalized beyond specific ladder configurations in structured environments (e.g., completely vertical ladders and no perturbations) [11]. For example, a humanoid equipped with dexterous hands and an a priori motion trajectory climbed a vertical ladder [12]. Further extensions of this work validated the robot on several different vertical ladders, but did not demonstrate improved speed (it was still confined to a slow pace of 26 mm/h vertical ascent) or robustness to perturbations [13]. In another work, a motion planner and a compliance controller were combined to generate disturbance-resistant climbing trajectories, but in the one ladder example demonstrated on hardware, the robot took seven minutes to traverse only five rungs [14].
Ladder climbing has also been demonstrated with a few quadrupeds [15, 16]. Like research to-date on humanoid ladder climbing, completely vertical ladders were the only subjects of these studies. Furthermore, the robots’ movements up the ladders were also slow, taking up to two minutes to ascend a single rung.
@@
Outside of ladder climbing specifically, robust locomotion of quadrupeds has been demonstrated in other challenging environments. Model-based methods, built around non-linear model predictive control or other optimization methods, typically excel in sparse terrains such as stepping stones and gaps [17]. However, such methods are vulnerable to modeling uncertainties, external disturbances, and degraded perception. In contrast, model-free methods such as reinforcement learning (RL) have shown great simulation to reality (sim2real) transfer, real-world robustness over rough terrain [18, 2, 19], and steady progress on sparse terrain problems such as stepping stones [20, 21, 22] and parkour [5]. Even so, no work to date has addressed the need for fast, robust, and generalizable ladder climbing with quadrupedal robots.
In this paper, we address long-standing challenges in robotic ladder climbing with the following contributions:
  • An RL framework for tracking position commands, enabling robust ladder climbing (Fig. 1; Supplementary Video).
@@
  • A hooked end-effector design that generates the necessary forces for secure and repeatable climbing (Fig. 2).
@@
  • Extensive simulations of successful ladder climbing across diverse configurations parameterized by $L_{len}$, $L_{width}$, $L_{space}$, $L_{radius}$, and $L_{\theta}$.
@@
  • Hardware demonstrations of the fastest and most generalized robotic ladder climbing to date, tested on ladders with varying $L_{\theta}$ and under unmodeled perturbations.
II Methods
@@
The general training process for the control policy is illustrated in Fig. 3 and follows that of reference [2]. First, in a simulation of the robot equipped with the hooked end-effectors, we train a teacher policy with access to noiseless proprioceptive observations $o_{p}$, noiseless inertial measurement unit (IMU) history $o_{i}^{H}$, a height scan around the robot $o_{e}$, and privileged state information $s_{p}$. The student policy is then distilled that only has access to noisy on-board proprioceptive measurements $\hat{o}_{p}$, IMU history $\hat{o}_{i}^{H}$, and ladder state $\hat{s}_{l}$. The resulting student policy outputs joint position targets $q_{i}$ at 50 Hz, which are tracked by a PD controller running at 400 Hz on the real robot. An neural network serves to model actuation dynamics in simulation [23].
@@
Figure 2: Quadruped robot used for testing, equipped with four hooked end-effectors. Composed of aluminum cores and 3D printed shells, the hook design features concave surfaces that encourage stable poses on round rungs, as well as protrusions that allow the robot to push and pull on the rungs.
@@
Figure 3: Illustration of the network architecture and training pipeline. First, a teacher policy is trained in simulation with access to noiseless observations and privileged state information. Then, a student policy with a recurrent belief encoder is trained to mimic the teacher actions and reconstruct the true privileged state from noisy observations and a noisy estimate of the ladder state and pose. The student policy is the one we deploy on the actual robot hardware.
II-A Policy Observations
@@
Proprioceptive observations $o_{p}$ include the commanded target position, target heading, gravity direction in base frame, joint positions, joint velocities, a history of joint tracking errors and velocities, a history of previous policy actions, and a binary flag indicating whether the agent is within 15 cm of the goal position. A history of the last eight IMU measurements $(o_{i}^{1},\dots,o_{i}^{8})$ at 400 Hz is also provided, which includes base linear acceleration and angular velocity. Direct IMU readings are used in lieu of the base velocities output by the ANYmal state estimator, as the latter was observed to produce linear errors up to 1 m/s while climbing.
A height scan around the robot $o_{e}$ with dimensions 2x1 m and a resolution of 10 cm is also input during teacher training because it helps the policy train faster. Later, $o_{e}$ is removed during student training, as regions above the robot are unobservable during deployment.
@@
Privileged state information $s_{p}$ includes body contact states, contact forces at the feet, feet friction coefficients, external forces and torques applied to the base, external forces applied to the feet, mass added to the base, airtime of each foot, feet positions in base frame, a vector of the ladder state $s_{l}$, and the ladder pose $p_{l}$. The ladder state comprises a binary flag that indicates whether a ladder is present in the current terrain, ladder inclination, width, the number of rungs, rung radius, and rung spacing. The ladder pose consists of the position and yaw of the bottom rung in the robot base frame. The noisy ladder state $\hat{s}_{l}$ and pose $\hat{p}_{l}$ are given as observations during student training and are estimated from motion capture data during real-world deployment.
II-B Teacher Policy Training
The teacher policy consists of four separate MLPs, each with elu activations (Fig. 3). Smaller MLPs first encode the height scan, privileged, and IMU observations before concatenating them with proprioceptive observations. A larger MLP then processes them into the final actions.
@@ II-B Teacher Policy Training
The teacher is trained using IPO [24] with adaptive constraint thresholding [19]. The rewards are summarized in Tab. III, and constraints are used to enforce the joint limits (position, velocity, torque). Episodes are terminated if the robot base inclination exceeds 100 in pitch or roll.
@@
We randomly generate two different types of terrain for training: i) rough terrain consisting of boxes and slopes, and ii) ladders of varying $L_{len}$, $L_{\theta}$, $L_{width}$, $L_{space}$, and $L_{radius}$. The training curriculum is adaptive, and agents progress to more difficult terrains (longer, steeper) as the agent reaches earlier goals. The rungs are parameterized as elliptic cylinders with a minor axis of 2.5 cm and a major axis that decreases with increasing curriculum difficulty, down to 2.5 cm. This value was chosen such that the rungs would be smaller than the 2.75 cm opening radius of the hook. The ladders are randomly offset from the end platform up to 15 cm, with a minimum clearance for the foot as $L_{\theta}$ increases.
@@
The agent is spawned in a random configuration and commanded to reach a random goal position and heading. For ladder terrains, the agent is always spawned at the base of the ladder, and the goal is always on the end platform. We randomly initialize the agent using a state from the previous episode half of the time, with additional randomization on the base and joint velocities. We find that this increases the range of states that the policy visits and improves the robustness of the final policy. Moreover, we do not sample command times, but use a dense tracking reward and allow any time up to the maximum episode length of 10 s. We find that this improves the success rate compared to prior methods [5] and allows the policy to learn emergent retry behaviors after slipping. We also apply random external forces and torques to the base and feet of the robot, and randomize the base mass and friction coefficients of the feet each episode.
@@
Heading Tracking $0.5\exp(10(\psi_{\mathrm{goal}}-\psi_{b})^{2})\exp(4||p_{\mathrm{goal}}||^{2})$
Base Motion $0.2(\exp(v_{b,z}^{2})+\exp(0.5(\dot{\phi}_{b}^{2}+\dot{\theta}_{b}^{2})))$
@@
Foot Slippage $-0.25\sum_{k\in\mathrm{feet}}c_{k}||v_{k}||(1-0.8\mathbf{I}(\mu_{k}<0.5))$
@@
Stand Still Contact $-0.5\delta_{\mathrm{goal}}\sum_{k\in\mathrm{feet}}\neg c_{j}$
@@
Base Collision $-c_{b}$
@@
$v_{b,\mathrm{over}}$ $(||v_{b}||-0.7)\cdot\mathbf{I}(v_{b}\cdot\hat{p}_{\mathrm{goal}}>0)$
@@
$\delta_{\mathrm{goal}}$ $\mathbf{I}(||p_{\mathrm{goal}}||<0.15)$
@@
$\mathbf{I}(\cdot)$ Identity function
II-C Student Policy Training
The student policy is trained in the same environment as the teacher, with identical initialization, domain randomization, and terrain curriculum. However, the student policy only has access to noisy versions of proprioceptive observations $\hat{o}_{p}$, IMU history $\hat{o}_{i}^{n}$, ladder state $\hat{s}_{l}$, and ladder pose $\hat{p}_{l}$. A small MLP encodes the noisy ladder state and pose before concatenating with the proprioceptive and latent IMU observations. This vector is passed to the belief encoder, which reconstructs the latent privileged state of the teacher, following the design of [2]. In this work, we treat the noisy ladder state and pose as exteroceptive observations. The latent privileged state is added back to the proprioceptive observations before passing to a larger MLP that produces the final actions.
The student is warm-started by copying over the matching network weights from the teacher. All other layers are randomly initialized. We collect rollouts using the student policy and compute the behavior loss by taking the mean squared error between the teacher and student actions [25]. The reconstruction loss is calculated via the L1 and L2 error between the predicted and actual privileged state. The network is trained using truncated backpropogation through time with a sequence length of 15.
A key advantage of privileged teacher-student training is that the teacher can be trained relatively quickly in the absence of noise and with knowledge of the applied disturbances. We found that the teacher struggled to learn an effective and robust policy when trained directly on student observations.
II-D Simulation Setup
@@
All simulations are performed in LeggedGym [26]. We train using 4096 parallel environments with 48 and 120 steps per batch for the teacher and student, respectively. The teacher is trained for 15,000 epochs and the student is trained for 5,000, taking around 4.5 days in total when trained on an RTX 3090. Training with the hook end-effector is around 30% slower than with the ball foot due to the additional collision bodies, which are approximated via convex decomposition.
III Results
III-A Simulation Results

### v2

@@ Abstract
Quadruped robots are proliferating in industrial environments where they carry sensor payloads and serve as autonomous inspection platforms. Despite the advantages of legged robots over their wheeled counterparts on rough and uneven terrain, they are still unable to reliably negotiate a ubiquitous feature of industrial infrastructure: ladders. Inability to traverse ladders prevents quadrupeds from inspecting dangerous locations, puts humans in harm’s way, and reduces industrial site productivity. In this paper, we learn quadrupedal ladder climbing via a reinforcement learning-based control policy and a complementary hooked end effector. We evaluate the robustness in simulation across different ladder inclinations, rung geometries, and inter-rung spacings. On hardware, we demonstrate zero-shot transfer with an overall 90% success rate at ladder angles ranging from 70 to 90, consistent climbing performance during unmodeled perturbations, and climbing speeds 232$\times$ faster than the state of the art. This work expands the scope of industrial quadruped robot applications beyond inspection on nominal terrains to challenging infrastructural features in the environment, highlighting synergies between robot morphology and control policy when performing complex skills. More information can be found at the project website: https://sites.google.com/leggedrobotics.com/climbingladders.
@@ I INTRODUCTION
Historically, most quadruped robots have been limited to locomotion on highly structured terrains [1]. Advances in control algorithms and hardware over the past decade have resulted in highly agile systems capable of stable locomotion over irregular natural surfaces [2, 3] and “parkour” over obstacles [4, 5]. Thanks to their ability to traverse rough terrain and remain stable on uneven surfaces, quadruped robots are becoming commonplace at industrial sites where they perform routine inspection tasks that are dangerous or undesirable for humans [6]. Despite ongoing advances, quadruped robots are still unable to robustly traverse many types of infrastructure common in man-made environments, a key example being ladders.
@@
Figure 1: Composite image of a quadrupedal robot equipped with hooked end effectors, ascending a ladder in 4 s with a reinforcement learning-based control policy. Ladder shown has parameters 90 incline, 1.8 m length, 0.3 m inter-rung spacing, 2.5 cm rung radius, and 1 m width.
@@
Ladder falls are a major occupational hazard; the American Academy of Orthopedic Surgeons estimates that 500,000 ladder-related injuries occur per year in the United States alone [7]. To minimize occupational risk and improve site efficiency, next-generation robots must be able to robustly negotiate all types of industrial terrain, including ladders. However, quadrupeds are not normally equipped with the appropriate morphology or control strategy for such a task. Consider that quadrupeds typically have ball-shaped or flat feet, preventing the anchoring forces needed for reliable upward movement [8]. Ladder climbing also requires full-body coordination to stabilize the center of mass and ascend steep inclines [9], a challenge for robots in unmodeled environments [10].
@@
Previous research on robotic ladder climbing has been conducted mainly with humanoid robots, is characterized by very slow climbing speeds, and has not generalized beyond specific ladder configurations in structured environments (e.g., completely vertical ladders and no perturbations) [11]. In one such work, a humanoid equipped with dexterous hands and an a priori motion trajectory climbed a vertical ladder [12]. Extensions to this work validated the robot on several different vertical ladders, but did not demonstrate improved vertical speed beyond $\sim$26 mm/h or robustness to perturbations [13]. In another work, a motion planner and compliance controller were combined to generate disturbance-resistant climbing trajectories. However, in the one ladder example demonstrated on hardware, the robot took seven minutes to traverse only five rungs [14]. Ladder climbing has also been demonstrated on a couple of quadrupeds [15, 16], although only vertical ladders were considered and the robots’ movements remained slow, taking up to two minutes to ascend a single rung.
@@
Outside of ladder climbing specifically, robust locomotion of quadrupeds has been demonstrated in other challenging environments. Model-based methods, built around non-linear model predictive control or other trajectory optimization methods, typically excel in sparse terrains such as stepping stones and gaps [17]. However, such methods are vulnerable to modeling uncertainties, external disturbances, and degraded perception. In contrast, model-free methods such as reinforcement learning (RL) have shown great simulation to reality (sim2real) transfer, real-world robustness over rough landscapes [2, 18], and steady progress on sparse terrain problems such as stepping stones [19, 20] and parkour [5].
In this paper, we propose a model-free RL control policy and a complementary hooked end effector that enables fast and generalizable quadrupedal ladder climbing (Fig. 1; Supplementary video). We extend [2] and use a privileged teacher-student training setup to first learn a robust teacher policy with access to noiseless information about the environment and external disturbances. We then distill this behavior into a student policy which only has access to on-board observations, but is trained to reconstruct the privileged information via a recurrent memory architecture.
Although elevation mapping is a common sensing modality for rough terrain locomotion [2, 18, 19], it fails in the case of vertical ladders (which appear as a line when viewed from above). In this paper, we focus on locomotion, so in lieu of more complex perception methods such as depth cameras [20] or terrain reconstruction [5], we use a motion capture system to provide the ladder position to the control policy. In summary, the contributions of this paper include:
  • An extension of [2] to generate robust control policies for quadrupedal ladder climbing with model-free RL.
@@
  • A hook end effector design that generates the necessary forces for reliable and robust climbing.
@@
  • Extensive simulations of successful climbing across diverse ladders parameterized by length $L_{\mathrm{len}}$, width $L_{\mathrm{width}}$, inter-rung spacing $L_{\mathrm{space}}$, rung radius $L_{\mathrm{radius}}$, and inclination angle $L_{\mathrm{\theta}}$.
@@
  • Hardware demonstrations of the fastest and most generalized robotic ladder climbing to date, tested on ladders with varying $L_{\mathrm{\theta}}$ and under unmodeled perturbations.
II Hook design
A hooked end effector was developed to encourage the emergence of robust climbing behavior. Multiple design variations were heuristically evaluated and iteratively refined through simulation (Fig. 2: Top). The final design incorporates strategically contoured concave regions, enabling the robot to exert both compressive and tensile forces on the rungs (Fig. 2: Bottom). This configuration facilitates stable postures, even when the robot’s center of mass extends beyond its support triangle. The hook’s large concave surface allows passively stable ladder engagement over a large range of angles, a feature that proved critical for maintaining performance under external disturbances.
III Control Methods
@@
The general training pipeline is illustrated in Fig. 3 and follows that of reference [2]. We first train a teacher policy in simulation with access to noiseless proprioceptive observations $o_{p}$, noiseless inertial measurement unit (IMU) history $o_{i}^{H}$, a height scan around the robot $o_{e}$, and privileged state information $s_{p}$. We then distill a student policy that only has access to noisy onboard proprioceptive measurements $\hat{o}_{p}$, IMU history $\hat{o}_{i}^{H}$, ladder state $\hat{s}_{l}$, and ladder pose $\hat{p}_{l}$. The resulting student policy outputs joint position targets $q_{i}$ at 50 Hz, which are tracked by a proportional-derivative controller running at 400 Hz on the real robot. A learned actuator network models the joint dynamics in simulation [21].
@@
Figure 2: Top: Different hook designs screened in simulation. Bottom: Quadruped robot used for testing, equipped with the best-performing hooked end effector. Composed of aluminum cores and 3D printed shells, the hook features concave surfaces that encourage stable poses on round rungs, as well as protrusions that allow pushing and pulling on the rungs. The robot’s base frame is defined with its origin at the center of the torso and the illustrated coordinate directions.
@@
Figure 3: Illustration of the network architecture and training pipeline. First, a teacher policy is trained in simulation with access to noiseless observations and privileged state information. Then, a student policy with a recurrent belief encoder is trained to mimic the teacher actions and reconstruct the true privileged state from noisy observations and a noisy estimate of the ladder state and pose. The student policy is deployed on the actual robot hardware.
@@
Policy Observations: Proprioceptive observations $o_{p}$ include the goal direction and heading in base frame, gravity direction, current joint positions, and a history of joint velocities and joint position tracking errors at 0 s, -0.02 s, and -0.04 s relative to the current timestep. A “standing mode” observation is provided which is true if the agent is within 15 cm of the goal, arbitrarily chosen. In lieu of base velocities, we directly provide a history of the last eight IMU measurements $(o_{i}^{1},\dots,o_{i}^{8})$ at 400 Hz, including base linear acceleration and angular velocity.
@@
During teacher training, a height scan $o_{e}$ with dimensions 2x1 m and a resolution of 10 cm is provided around the robot, as it was found to accelerate policy learning. It is subsequently removed during student training. Privileged state information $s_{p}$ includes body contact states, feet contact forces, friction coefficients, external forces and torques applied to the base and feet, mass added to the base, airtime of each foot, feet positions in base frame, a ladder state vector $s_{l}$, and ladder pose $p_{l}$. The ladder state comprises a binary flag indicating whether a ladder is present in the current terrain, $L_{\mathrm{\theta}}$, $L_{\mathrm{width}}$, $L_{\mathrm{radius}}$, $L_{\mathrm{space}}$, and number of rungs. The ladder pose consists of the position and yaw of the bottom rung in the robot base frame. A noisy version of the ladder state $\hat{s}_{l}$ and pose $\hat{p}_{l}$ are given as exteroceptive observations during student training. During real-world deployment, the ladder state is measured directly, and the pose is estimated from motion capture data. Noise is added to the state in simulation to account for measurement error and irregularities in ladder geometry.
@@
Teacher Policy Training: The teacher policy is similar to that of [2], with the addition of a multi-layer perceptron that encodes the IMU observation history before concatenating it with proprioceptive observations. The policy is trained using Interior-point Policy Optimization (IPO) with adaptive constraint thresholding [18]. IPO is used to enforce constraints on the joint limits (position, velocity, torque), which was found to result in fewer violations compared to a reward penalty. Tab. III summarizes the rewards. Episodes terminate if the robot base inclination exceeds 100 in pitch or roll.
@@
Two different types of terrain are randomly generated during training: i) rough terrain consisting of boxes and slopes, and ii) ladders of varying $L_{\mathrm{len}}$, $L_{\mathrm{\theta}}$, $L_{\mathrm{width}}$, $L_{\mathrm{space}}$, and $L_{\mathrm{radius}}$. The training curriculum is adaptive, and agents progress to more difficult terrains (longer, steeper) as the agent reaches earlier goals [22]. To simplify the curriculum, the rungs are parameterized as elliptic cylinders with a minor axis of 2.5 cm and a major axis that decreases with increasing curriculum difficulty, down to 2.5 cm. This value was chosen such that the rungs would be smaller than the 2.75 cm opening radius of the hook. Early-curriculum ladders therefore have broad, flat rungs that are easy to climb. The ladders are randomly offset from the end platform up to 15 cm, with a minimum clearance for the foot as $L_{\theta}$ increases.
@@
Agents are spawned in a random configuration and commanded to reach a random goal position and heading. Unlike [5], we do not sample command times, as we find that i) it is difficult to estimate the amount of time required for climbing, and ii) a fixed command time prevents the agent from falling and learning robust retry behaviors. Instead, we use a dense tracking reward (see Tab. I) and allow any time up to the episode length of 10 s. We find that this improves success rates compared to prior methods. The agent is initialized using a state from the previous episode $50\%$ of the time, with randomization added to the base and joint velocities. We also apply random external forces and torques to the base of the robot, random velocity offsets (pushes), randomize the base mass, and randomize the friction coefficients of the feet.
@@
Heading Tracking $0.5\exp(-10(\psi_{\mathrm{goal}}-\psi_{b})^{2})\exp(-4||p_{\mathrm{goal}}||_{2}^{2})$
Base Motion $0.2(\exp(-v_{b,z}^{2})+\exp(-0.5(\dot{\phi}_{b}^{2}+\dot{\theta}_{b}^{2})))$
@@
Foot Slippage $-0.25\sum_{k\in\mathrm{feet}}c_{k}||v_{k}||_{2}(1-0.8\mathbf{1}(\mu_{k}<0.5))$
@@
Stand Still Contact $-0.5\delta_{\mathrm{goal}}\sum_{k\in\mathrm{feet}}\neg c_{k}$
@@
Base Collision $-c_{\mathrm{base}}$
@@
$v_{b,\mathrm{over}}$ $\max(0,||v_{b}||_{2}-0.7)$
@@
$\delta_{\mathrm{goal}}$ $\mathbf{1}(||p_{\mathrm{goal}}||_{2}<0.15)$
@@
$\mathbf{1}(\cdot)$ Binary indicator function
Student Policy Training: A student policy is trained using supervised learning to output the same actions as the teacher policy, while only having access to noisy versions of the observations. A recurrent belief encoder is used to reconstruct the privileged observations provided to the teacher. In this work, the noisy ladder state and pose are treated as exteroceptive observations that the student can reconstruct from contacts with the environment.
@@
Simulation Setup: All simulations are performed in LeggedGym [22]. We train using 4096 parallel environments with 48 and 120 steps per batch for the teacher and student, respectively. The teacher is trained for 15,000 epochs and the student is trained for 5,000, taking around 4.5 days in total on a NVIDIA RTX 3090. Training with the hook end effector is around 30% slower than with the ball foot due to the additional collision bodies, which are approximated by convex decomposition.
IV Results